# AnyDressing: Customizable Multi-Garment Virtual Dressing via Latent Diffusion Models

### v1

@@ 1 Introduction
VD is used to be regarded as a subtask of traditional subject-driven image customization, prior approaches [10, 18, 25, 30, 39, 41, 45, 57] simply integrate the features of reference image into the text embeddings without fully exploiting the information from the reference image. Several subsequent works [54, 32] more comprehensively utilize the features of the reference image by training additional cross-attention layers to integrate reference image features into the diffusion model. However, these methods struggle to preserve the intricate textures of the garment. Recently, some methods [4, 47, 40] focus on garment-centric image generation. Most of them leverage a full copy of diffusion U-Net as the garment encoder named ReferenceNet to maintain fine-grained garment information. Nevertheless, these methods are tailored exclusively to single items of clothing and lack support for multiple conditions, thus hindering the ability to freely dress in any combination of various garments.
@@
To address the aforementioned issues, we propose AnyDressing, a novel approach that customizes characters conditioned on any combination of garments and any personalized text prompts. AnyDressing primarily comprises two primary networks named GarmentsNet and DressingNet. The GarmentsNet leverages a core Garment-Specific Feature Extractor (GFE) module to extract multi-garment detailed features, which utilizes parallelized self-attention LoRA layers within a shared U-Net architecture to individually encode garment textures. This design not only avoids clothing blending but also ensures network efficiency, allowing for easy scalability to any number of garments. The DressingNet employs a Dressing-Attention (DA) mechanism to seamlessly integrate multi-garment features into the denoising process. To ensure that each garment instance focuses specifically on its corresponding region, we further introduce a novel Instance-Level Garment Localization (IGL) learning strategy in DA. This avoids influencing other irrelevant regions in the synthetic image, thus improving fidelity to arbitrary customized text prompts. Additionally, to enhance texture details, we design a Garment-Enhanced Texture Learning (GTL) strategy that strengthens the supervision of attire by imposing constraints from perceptual features and high-frequency information.
@@ 2 Related Work
Virtual Dressing. Virtual Dressing (VD) [4, 47, 40] aims to generate freely editable human images with reference garments and optional conditions. StableGarment [47] and IMAGDressing [40] leverage a garment U-Net for extracting fine-grained clothing features and a denoising U-Net with a hybrid attention module to incorporate garment features into denoising process. Magic Clothing [4] additionally proposes a joint classifier-free guidance to balance the control of garment features and text prompts. However, existing approaches are limited to processing single items of clothing, and difficult to maintain fidelity to text prompts. In contrast, our method allows for freely dressing multiple garments and produces coherent and attractive images following customized text prompts.
@@ 4 Methodology
Given $N$ target garments, the proposed AnyDressing aims to generate a new image $x_{dr}$, showcasing a customized character dressed in multiple target garments across a variety of scenes, styles and actions based on the text prompt. As illustrated in Fig. 2, AnyDressing comprises two primary networks: GarmentsNet and DressingNet. The GarmentsNet leverages the Garment-Specific Feature Extractor module to extract detailed features from multiple garments (Sec. 4.1). Meanwhile, the DressingNet integrates these features for virtual dressing using a Dressing-Attention module and an Instance-Level Garment Localization Learning mechanism (Sec. 4.2). Additionally, a Garment-Enhanced Texture Learning strategy is designed to further enhance crucial texture details in the synthesis images (Sec. 4.3). Next, we will introduce the aforementioned modules, along with training and inference processes (Sec. 4.4), in detail.
@@ 4.1 GarmentsNet
Drawing inspiration from the successful practice of the aforementioned reference mechanisms, we observe that self-attention layers are crucial for the implicit warping of features, enabling the effective matching of input garments to the appropriate body parts. Meanwhile, other layers are typically responsible for general feature extraction and can be shared across different garments without compromising the model’s performance. Building on this insight, we innovatively design a simple yet effective architecture named GarmentsNet, which employs a core Garment-Specific Feature Extractor (GFE) module to individually encode features for each garment utilizing LoRA [15] layers within a shared U-Net framework. As a result, this design significantly avoids garment blending while maintaining network efficiency. As illustrated in Fig. 2, the GFE module employs a parallelized self-attention mechanism to extract detailed features of multiple garments. Specifically, for each garment condition $\mathbf{F}_{i}$, we define the proprietary self-attention LoRA matrix $\triangle\hat{\mathbf{W}}_{i}$ as follows:
@@
where $\mathbf{Q}_{i}=\mathbf{F}_{i}(\hat{\mathbf{W}}_{q}+\triangle\hat{\mathbf{W}}_{q}^{i})$, $\mathbf{K}_{i}=\mathbf{F}_{i}(\hat{\mathbf{W}}_{k}+\triangle\hat{\mathbf{W}}_{k}^{i})$, $\mathbf{V}_{i}=\mathbf{F}_{i}(\hat{\mathbf{W}}_{v}+\triangle\hat{\mathbf{V}}_{v}^{i})$, only $\triangle\hat{\mathbf{W}}$ is trainable and $N$ represents the number of reference garments.
Thanks to the multi-garment parallel processing design of our GFE module, GarmentsNet can seamlessly scale to any number of garments. Notably, this expansion requires only some newly added $\triangle\hat{\mathbf{W}}$, and significantly minimizes both training and inference time compared with duplicating the entire garment encoding network.
@@ 4.2.1 Adaptive Dressing-Attention
In the VD task, the main denoising network is typically kept frozen during training [4, 40] to preserve its original editing and generation capabilities as much as possible. However, this prevents the pre trained attention modules from effectively capturing the relationship between reference garment features and latent features. To address this issue, we design an adaptive Dressing-Attention (DA) mechanism to efficiently integrate multi-garment texture cues into synthetic images, inspired by [54]. As shown in Fig. 2, the Dressing-Attention module includes a frozen self-attention module and a learnable cross-attention module. Let $\{\mathbf{F}_{1},\mathbf{F}_{2},\cdots,\mathbf{F}_{N}\}$ denote $N$ garment features output by the GarmentsNet at corresponding positions, we first concatenate these features along the spatial dimension to obtain the final garment features: $\mathbf{F}_{all}=Concat(\mathbf{F}_{1},\mathbf{F}_{2},\cdots,\mathbf{F}_{N})$. We then introduce two trainable linear projection layers $\mathbf{W}_{k}^{\prime}$ and $\mathbf{W}_{v}^{\prime}$ to align garment features with latent feature $\mathbf{Z}$. Formally, the output of Dressing-Attention $\mathbf{Z}_{new}$ is:
@@ 5.1 Setup
Notably, a dataset comprising image triplets that include model images paired with multiple laid-out garments is currently lacking. Therefore, we utilize a HumanParsing model to extract clothing items from DressCode [33] and an additional proprietary dataset collected from the internet, forming triplet data pairs (upper garment, lower garment, person image). In these triplets, one garment is an original laid-out image, while the other is a segmented image from the person’s image. Finally, we construct 26,114 public triplets from Dresscode and 37,065 triplets from proprietary dataset to train AnyDressing. For model evaluation, we introduce two benchmarks to evaluate the model on single-garment and multi-garment dressing respectively. Specifically, for single-garment evaluation, we handpick 300 diverse garments from VITON-HD [5] encompassing various styles and colors, and additionally gather 300 reference garments with intricate textures and non-garment images from the internet. For multi-garment evaluation, we meticulously collect 25 lowers from the internet and pair each with 10 different uppers, resulting in a total of 250 pairs, called Dressing-Pair. We generate images for each test garment with the provided 7 text prompts.
Implementation Details. In our experiments, we initialize the weights of GarmentsNet and DressingNet with the weights of Stable Diffusion v1.5 [38] and we use CLIP ViT-L/14 [36] as the text encoder. Our model is trained on paired images with a resolution of $768\times 576$. The model is trained on 8 NVIDIA A100 GPUs for 100k steps with a batch size of 4. We adopt AdamW [31] optimizer with a fixed learning rate of 5e-5. During inference, we use DDIM [43] sampler with 30 steps and set guidance scale $\omega$ to 6.0. For more detailed settings, please refer to the supplementary materials.
Baselines. We compare our method against the following state-of-the-art image synthesis method: IP-Adapter [54], MagicClothing [4], StableGarment [47] and IMAGDressing [40]. We directly use their released pre-trained models for comparison. For a fair comparison, all experiments are conducted with a resolution of $768\times 576$.
Evaluation Metrics. Similar to previous methods, we quantitatively evaluate the generated results from three perspectives: CLIP-T for text consistency, CLIP-I for texture consistency, and Aesthetic Score (AS) for overall generation quality. Especially, to better evaluate multi-garment dressing, we introduce a new metric CLIP-$\text{I}^{*}$ to assess texture consistency by leveraging OpenPose [1] to obtain the matching partitions of the reference garments in the synthesized image and averaging their CLIP-I metrics.
@@ 5.4 Ablation Studies
Figure 5: Ablation results on GFE and IGL modules.
Figure 6: Ablation results on GTL module.
@@ 7.1 Detailed Parameters
In our experiments, we use SOTA large multi-modal model CogVLM [48] to caption the image. The hyper-parameters used in our experiments are set as follows:

### v2

@@ 1 Introduction
VD is used to be regarded as a subtask of traditional subject-driven image customization, prior approaches [10, 18, 25, 31, 39, 41, 45, 57] simply integrate the features of reference image into the text embeddings without fully exploiting the information from the reference image. Several subsequent works [54, 33] more comprehensively utilize the features of the reference image by training additional cross-attention layers to integrate reference image features into the diffusion model. However, these methods struggle to preserve the intricate textures of the garment. Recently, some methods [4, 47, 40] focus on garment-centric image generation. Most of them leverage a full copy of diffusion U-Net as the garment encoder named ReferenceNet to maintain fine-grained garment information. DreamFit [30] proposes a lightweight garment encoder, which utilizes trainable LoRA layers to extract garment features instead of finetuning a full copy of the UNet. Nevertheless, these methods are tailored exclusively to single items of clothing and lack support for multiple conditions, thus hindering the ability to freely dress in any combination of various garments.
@@
To address the aforementioned issues, we propose AnyDressing, a novel approach that customizes characters conditioned on any combination of garments and any personalized text prompts. AnyDressing primarily comprises two primary networks named GarmentsNet and DressingNet. The GarmentsNet leverages a core Garment-Specific Feature Extractor (GFE) module to extract multi-garment detailed features, which utilizes parallelized self-attention layers within a shared U-Net architecture to individually encode garment textures. And we employ LoRA mechanism within the self-attention layers to further reduce the parameter increase associated with the added garments. The GFE module not only avoids clothing blending but also ensures network efficiency, allowing for easy scalability to any number of garments. The DressingNet employs a Dressing-Attention (DA) mechanism to seamlessly integrate multi-garment features into the denoising process. To ensure that each garment instance focuses specifically on its corresponding region, we further introduce a novel Instance-Level Garment Localization (IGL) learning strategy in DA. This avoids influencing other irrelevant regions in the synthetic image, thus improving fidelity to arbitrary customized text prompts. Additionally, to enhance texture details, we design a Garment-Enhanced Texture Learning (GTL) strategy that strengthens the supervision of attire by imposing constraints from perceptual features and high-frequency information.
@@ 2 Related Work
Virtual Dressing. Virtual Dressing (VD) [4, 47, 40] aims to generate freely editable human images with reference garments and optional conditions. StableGarment [47] and IMAGDressing [40] leverage a garment U-Net for extracting fine-grained clothing features and a denoising U-Net with a hybrid attention module to incorporate garment features into denoising process. Magic Clothing [4] additionally proposes a joint classifier-free guidance to balance the control of garment features and text prompts. DreamFit [30] proposes a lightweight garment encoder based on trainable LoRA layers to streamline model complexity and memory usage. However, existing approaches are limited to processing single items of clothing, and difficult to maintain fidelity to text prompts. In contrast, our method allows for freely dressing multiple garments and produces coherent and attractive images following customized text prompts.
@@ 4 Methodology
Given $N$ target garments, the proposed AnyDressing aims to generate a new image $x_{dr}$, showcasing a customized character dressed in multiple target garments across various scenes, styles and actions based on the text prompt. As illustrated in Fig. 2, AnyDressing comprises two primary networks: GarmentsNet and DressingNet. The GarmentsNet leverages the Garment-Specific Feature Extractor module to extract detailed features from multiple garments (Sec. 4.1). Meanwhile, the DressingNet integrates these features for virtual dressing using a Dressing-Attention module and an Instance-Level Garment Localization Learning mechanism (Sec. 4.2). Additionally, a Garment-Enhanced Texture Learning strategy is designed further to enhance crucial texture details in the synthesis images (Sec. 4.3). Next, we will introduce the aforementioned modules, along with training and inference processes (Sec. 4.4), in detail.
@@ 4.1 GarmentsNet
Drawing inspiration from the successful practice of the aforementioned reference mechanisms, we observe that self-attention layers are crucial for the implicit warping of features, enabling the effective matching of input garments to the appropriate body parts. Meanwhile, other layers are typically responsible for general feature extraction and can be shared across different garments without compromising the model’s performance. Building on this insight, we innovatively design a simple yet effective architecture named GarmentsNet, which employs a core Garment-Specific Feature Extractor (GFE) module to encode features for each garment utilizing individual self-attention layers within a shared U-Net framework. Inspired by [30], we integrate LoRA [15] mechanism into self-attention layers, minimizing the increase in parameters associated with the added garments. As a result, this design significantly avoids garment blending while maintaining network efficiency. As illustrated in Fig. 2, the GFE module employs a parallelized self-attention mechanism to extract detailed features of multiple garments. Specifically, for each garment condition $\mathbf{F}_{i}$, we define the proprietary self-attention LoRA matrix $\triangle\hat{\mathbf{W}}_{i}$ as follows:
@@
where $\mathbf{Q}_{i}=\mathbf{F}_{i}(\hat{\mathbf{W}}_{q}+\triangle\hat{\mathbf{W}}_{q}^{i})$, $\mathbf{K}_{i}=\mathbf{F}_{i}(\hat{\mathbf{W}}_{k}+\triangle\hat{\mathbf{W}}_{k}^{i})$, $\mathbf{V}_{i}=\mathbf{F}_{i}(\hat{\mathbf{W}}_{v}+\triangle\hat{\mathbf{W}}_{v}^{i})$, only $\triangle\hat{\mathbf{W}}$ is trainable and $N$ represents the number of reference garments.
Thanks to the multi-garment parallel processing design of our GFE module, GarmentsNet can seamlessly scale to any number of garments. Notably, this expansion requires only some newly added LoRA matrix $\triangle\hat{\mathbf{W}}$ in self-attention layers, and significantly minimizes both training and inference time compared with duplicating the entire garment encoding network. Considering the capability of the GFE module to individually encode each garment, we excise the cross-attention modules in GarmentsNet to further reduce redundancy.
@@ 4.2.1 Adaptive Dressing-Attention
In the VD task, the main denoising network is typically kept frozen during training [4, 40] to preserve its original editing and generation capabilities as much as possible. To incorporate reference garment features into latent features, we design an adaptive Dressing-Attention (DA) mechanism to efficiently integrate multi-garment texture cues into synthetic images, inspired by [54]. As shown in Fig. 2, the Dressing-Attention module includes a frozen self-attention module and a learnable cross-attention module. Let $\{\mathbf{F}_{1},\mathbf{F}_{2},\cdots,\mathbf{F}_{N}\}$ denote $N$ garment features output by the GarmentsNet at corresponding positions, we first concatenate these features along the spatial dimension to obtain the final garment features: $\mathbf{F}_{all}=Concat(\mathbf{F}_{1},\mathbf{F}_{2},\cdots,\mathbf{F}_{N})$. We then introduce two trainable linear projection layers $\mathbf{W}_{k}^{\prime}$ and $\mathbf{W}_{v}^{\prime}$ to align garment features with latent feature $\mathbf{Z}$. Formally, the output of Dressing-Attention $\mathbf{Z}_{new}$ is:
@@ 5.1 Setup
Notably, a dataset comprising image triplets that include model images paired with multiple laid-out garments is currently lacking. Therefore, we utilize a HumanParsing model to extract clothing items from DressCode [34] and an additional proprietary dataset collected from the internet, forming triplet data pairs (upper garment, lower garment, person image). In these triplets, one garment is an original laid-out image, while the other is a segmented image from the person’s image. Finally, we construct 26,114 public triplets from Dresscode and 37,065 triplets from proprietary dataset to train AnyDressing. For model evaluation, we introduce two benchmarks to evaluate the model on single-garment and multi-garment dressing respectively. Specifically, for single-garment evaluation, we select 300 reference garments from VITON-HD [5] encompassing various styles and colors, and additionally collect 300 diverse garments with intricate textures from the internet. For multi-garment evaluation, we meticulously gather 25 lowers from the internet and pair each with 10 different uppers, resulting in a total of 250 pairs, called Dressing-Pair. We generate images for each test garment with the provided 7 text prompts.
@@
Implementation Details. In our experiments, we initialize the weights of GarmentsNet and DressingNet with the weights of the U-Net in Stable Diffusion v1.5 [38]. Our model is trained on paired images at the resolution of $768\times 576$. The trainable parameters are GarmentsNet and the cross-attention layers in Dressing-Attention module. During training, We adopt AdamW [32] optimizer with a fixed learning rate of 5e-5. The model is trained for 100k steps on 8 NVIDIA A100 GPUs with a batch size of 4. During inference, we use DDIM [43] sampler with 30 steps and set guidance scale $\omega$ to 6.0. Please refer to the supplementary materials for more details.
Baselines. We compare our method against the following state-of-the-art image synthesis method: IP-Adapter [54], MagicClothing [4], StableGarment [47] and IMAGDressing [40]. We use the official model parameters from their official implementations. For a fair comparison, all experiments are conducted with the resolution of $768\times 576$.
Evaluation Metrics. We follow previous methods to adopt three metrics for evaluation: CLIP-T for text-image similarity, CLIP-I for garment consistency, and CLIP Aesthetic Score (CLIP-AS) for overall generation quality. Especially, to better evaluate multi-garment dressing, we introduce a new metric CLIP-$\text{I}^{*}$ to assess texture consistency by leveraging OpenPose [1] to obtain the matching partitions of the reference garments in the synthesized image and averaging their CLIP-I metrics.
@@ 5.3 Quantitative Comparisons
Figure 5: Ablation results on GFE and IGL modules.
Figure 6: Ablation results on GTL module.
@@ 7.1 Detailed Parameters
In our experiments, we use SOTA large multi-modal model CogVLM [48] to caption the image. GarmentsNet requires only one step forward process before the multiple denoising steps in DressingNet, causing a minimal amount of extra computational cost. The hyper-parameters used in our experiments are set as follows: